# Holistic Optimization Framework for FPGA Accelerators

### v3

@@ 15 G[i][j] += E[i][k] * F[k][j];} // S5
Listing 1: Code of 3mm
Dataflow  The dataflow method creates six tasks, each corresponding to a statement, with communication managed through FIFOs. For the 3mm kernel, efficient data reuse is critical and can be achieved via reuse buffers or repeated off-chip data transfers. However, reuse buffers are limited to individual tasks and cannot be shared across tasks.
As 3mm is computation-bound, increasing parallelism is vital. However, parallelism is constrained by the receiving task’s FIFO bit width, and attempts to scale this approach often lead to routing congestion and impractical bitstream generation. Despite these challenges, dataflow allows concurrent execution of the first two matrix multiplications and starts the third as soon as data becomes available.
Mixed Method  Methods like Allo [14] and HIDA [78] integrate shared buffers and dataflow techniques to enhance QoR. Allo extends HeteroCL [38] within the MLIR framework [40], supporting code transformation and dataflow but lacks automatic data-tile selection and pragma insertion.
Limitations in General  Existing HLS frameworks often lack key optimizations, such as automatic ping-pong buffering for overlapping computation and communication, tile size adjustments, and padding to maximize bit width for reducing communication latency. These gaps can lead to suboptimal performance, particularly in data-intensive scenarios. Most frameworks [60, 59] restrict unroll factors to evenly divide problem sizes, while others [67, 66] leveraging Merlin [77] permit uneven unroll factors. However, optimizations typically target complete tiles, resulting in inefficiencies with partial tiles and non-innermost unrolled loops. Expanding unroll factor options could better balance parallelism and performance.
For the 3mm kernel optimized with RTL simulation, our framework (Prometheus) achieved a throughput of 368.36 GF/s, significantly surpassing Sisyphus (178.97), Allo (60.40), ScaleHLS (43.04), and AutoDSE (1.74). These frameworks employ varying strategies—AutoDSE, Sisyphus, and ScaleHLS use shared buffer methods, while Allo applies a mixed method. However, due to the limitations noted, further improvements are needed to achieve a superior QoR.
II-C Overview of Prometheus
We propose Prometheus, a unified optimization framework leveraging a Non-Linear Problem (NLP) cost model to explore design spaces efficiently. Prometheus is engineer-friendly, offering adjustable parameters such as tile size, array partitioning, and unroll factors, while being SLR-aware for task distribution across SLRs, simplifying bitstream generation and optimizing performance-resource balance.
The pseudo-code in Listing 2 illustrates how Prometheus processes the 3mm kernel. The load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFOs. Each task_i represents the fully unrolled computation of the intra-tile for statement Si in Listing 1. We can see an example on Listing 3. Arrays E, F, and G are initialized to 0 with statements S0, S2, and S4, respectively, and are not loaded. Prometheus optimizes computation by fusing statements producing the same outputs (e.g., Listing 1), automatically generating an NLP problem to determine loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding. It strategically adjusts bit widths (e.g., for arrays F, D, and G) to balance parallelism and resource usage.
@@ II-C Overview of Prometheus
7 task_0(E); // S0 + intra-tile loops
@@
10 task_1(E, A, B); // S1 + intra-tile loops
@@
18 task_2(F); // S2 + intra-tile loops
@@
21 task_3(F, C, D); // S3 + intra-tile loops
@@
31 task_4(G); // S4 + intra-tile loops
@@
34 task_5(G, E, F); // S5 + intra-tile loops
@@
Listing 2: Pseudo Code of 3mm generated by Prometheus
Efficient overlap of computation and communication is achieved via ping-pong buffering, concurrent task execution, and FIFO-triggered dependent tasks. Fused tasks 0 and 1 execute concurrently, and Task 2 starts once arrays F and E are ready. These optimizations enable Prometheus to surpass frameworks like Sisyphus [60] in throughput and QoR, particularly for computation-bound kernels. Its SLR-awareness and dataflow integration make it practical for generating deployable, resource-efficient bitstreams.
III Problem Formulation
In our dataflow model, we adopt a synchronous dataflow where the sizes of the arrays are known during compile time. This compile-time awareness enables us to construct a precise model that facilitates rigorous optimizations. To leverage FPGA parallelism, we implement an acyclic dataflow graph, ensuring parent nodes do not receive data from their children. While this constraint limits graph configurations and may increase resource usage, it reduces overall latency. To support this structure, we inline [15] each function in the input code to generate the required acyclic graph. Our primary objective is to minimize latency within resource constraints by overlapping communication and computation within tasks and executing independent tasks concurrently. To achieve this, we apply various transformations and optimizations, explored in this section, and navigate the design space using an NLP-based approach, detailed in Section IV.
Dependency Graph Creation  Our process begins with the input C/C++ code, which undergoes maximal distribution to ensure each loop body contains only one statement, provided no dependencies exist within the loop. ISCC [69] verifies the legality of these transformations, ensuring dependencies are preserved. After achieving full distribution, where each loop body corresponds to a single task, we construct a dependency graph capturing inter-task dependencies. Tasks with identical outputs are then merged (when legal), creating fused tasks with output-stationary properties. If a dependency prevents distribution, the framework will still function but with a more limited optimization space, making it more effective when distribution is possible.
@@ III Problem Formulation
Data-tiling and Padding  From the fused dependency graph, edges represent data communication between tasks and off-chip memory, transferring data-tiles of specified sizes. Data tiling divides loop iterations into smaller tiles, splitting array accesses into subsets. Each loop $l$ iterating over an array $a$ is divided into an outer loop ($TC^{l}_{inter}$) and an inner loop ($TC^{l}_{intra}$), with feasible permutations applied. If an array is not fully transferred in a dimension $n$, its tile size is constrained by other arrays sharing the same loop. To optimize memory transfers, padding is applied to arrays in two ways. Simple padding increases the bit width ($BW_{a}$) for efficient transfers while maintaining the original loop trip count. Composite padding adjusts both $BW_{a}$ and the loop trip count ($TC^{l}$) to support unroll factors that do not evenly divide the original trip count, allowing irregular tile sizes and expanding the design space. Tile sizes are consistent within a fused task but vary between tasks. For example, in Listing 2, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
@@
Fine-grained Parallelism  Data-tile selection involves splitting each loop and permuting the outermost loop. Using ISCC [69], we ensure the legality of permutations, yielding two loop levels: inter-tile (outer) and intra-tile (inner). If permutation is infeasible, the inter-tile loop remains in its legal position. In Listing 2, inter-tile loops (lines 3 and 5) iterate intra-tile loops in task_0, which are fully unrolled to maximize parallelism. Array partitioning, determined by the unroll factor, ensures data resides in separate BRAM banks for simultaneous access. Reduction loops across inter-tiles are pipelined with an initiation interval ($II=n>1$), where n matches the reduction latency. For instance, in Listing 2, additions take 3 cycles, resulting in $II=3$. Fused tasks are output-stationary, ensuring each tile’s output is processed (loaded, computed, and stored or sent) only once. Dependent tasks execute sequentially with high throughput, leveraging unrolled tasks and on-chip data. Processed data-tiles are then transferred to the next task or off-chip memory, optimizing performance.
@@
Loop Order  Due to the full unrolling of the intra-task, there is no need to select the loop order within the intra-tile. We place the inter-tile reduction loops directly above the task and are pipelined. If multiple reduction loops exist, we rank them by the size of their trip counts, placing the loop with the highest trip count innermost to ensure an efficient pipeline. This setup provides the flexibility to choose the order of the inter-tile loops that are not reduction loops. This order will subsequently be determined by the NLP, as described in Section IV. As shown in Listing 2, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 1.
@@
Automatic Overlapping of Communication and Computation  To overlap communication and computation, we use on-chip buffers. The buffer size and data transfer location are determined by various options explored via NLP (cf. Section IV). Since data must stay on-chip for intra-task computation, transfers occur either below an inter-tile loop or before any loops start. The transfer position determines the data tile size, covering all data accessed below it. To improve data reuse, buffer size can match or exceed the transferred data tile. Similar to transfer location, buffer size is determined by its position relative to inter-tile loops or before any loops. Two boolean variables, $d_{a,l}$ and $t_{a,l}$, define and transfer array $a$ under loop $l$ when set to true. If defined or transferred before loops, $l=0$. Double-buffering is used for read-only arrays, and triple-buffering for arrays that are both read and written. Following [18], we perform an initial load, then overlap loading the next tile with computing the current one.
@@
Coarse-Grained Parallelism: Automatic Execution of Concurrent Tasks  Due to the use of the dataflow pragma, tasks can begin computation as soon as they have sufficient data, allowing for concurrent execution. This concurrency significantly increases overall performance.
@@
Memory Transfer  Communication latency is reduced by transferring more data per cycle using increased bit width with padding. Read-only arrays are duplicated in off-chip memory for tasks with multiple reads, eliminating feed-through logic. For non-read-only arrays, data passes between tasks via FIFOs, using the same bit width as for off-chip memory communication.
IV NLP Formulation
@@
To determine the tile size, loop order, bit width, and memory transfers, we formulate a cost model as a NLP problem aimed at minimizing overall latency. This approach builds upon the methodology proposed by Pouget et al. [59, 58, 60], which we have adapted to address our specific requirements and constraints. In our work, we incorporate dataflow considerations along with all the optimizations detailed in Section III. We employ PoCC [54] to extract compile-time information such as schedules, loop trip counts, dependencies, and operation counts per statement. ISCC [69] then generates all legal permutations for each loop body.
Constant Description
@@
$p_{i}^{l}$ Position of the permutation $i$ for the loop $l$
@@
TABLE I: Overview of the constant, variable and set employed in formulating the NLP
IV-A Constraints
We now describe the constraints by using the code of Listings 1, 2 and 3.

### v4

@@ 15 G[i][j] += E[i][k] * F[k][j];} // S5
Listing 3: Reference Implementation of the 3mm Kernel from PolyBench. This listing shows the baseline version of the 3mm benchmark, comprising three matrix multiplication operations. The first two multiplications are independent and produce intermediate results that serve as inputs to the final matrix multiplication.
2.4. Overview of Prometheus
We introduce Prometheus, a unified optimization framework designed to efficiently navigate complex design spaces using a Non-Linear Programming (NLP) cost model. Prometheus streamlines FPGA design development by providing an intuitive, engineer-friendly interface with tunable parameters, including tile size, array partitioning, and loop unrolling factors.
The framework integrates all the optimization techniques discussed in previous sections, ensuring a comprehensive and cohesive approach to performance enhancement. Additionally, Prometheus is SLR-aware, enabling intelligent task distribution across Super Logic Regions (SLRs). This capability not only simplifies bitstream generation but also optimizes the trade-off between computational performance and resource utilization.
Prometheus C++ Code Intermediate Representation (dependencies, trip counts) Task-flow Graph
1234
Formulate a Non-Linear Problem encoding the full optimization design space All Parameters of the Spacee.g. data-tile of array $A$ is $\texttt{32}\times\texttt{64}$Files for HLS CompilerHLS-C++, OpenCL, Config Bitstream with HLS Compiler
Extract with PoCC GenerateSolve NLP Generate
Figure 1. Overview of the Prometheus Framework Workflow. This diagram illustrates the end-to-end pipeline of the Prometheus optimization framework, starting from C++ source code and proceeding through intermediate representation extraction, task-flow graph generation, Non-Linear Programming (NLP)-based design space exploration, and final compilation into FPGA bitstreams using HLS compilers.
Group Parameter Description
Design Variables
Memory Bit Width ($BW_{a}$) Width of memory transfer in bits;
Data-tile Size (modeled via Transfer Level $t_{a,l}$ in Section 4) Size of the data tile of array $a$ transferred either from off-chip to on-chip memory or between tasks. The transfer level determines this size.
Data-tile Reuse Size (modeled via Reuse Level $d_{a,l}$ in Section 4) Size of the data tile of array $a$ reused within a task. The reuse level dictates this tile size.
Buffering ($N_{a}$) Number of buffers (2 for double buffering, 3 if read & write), enabling overlap of load/compute/store.
Communication Padding Additional data added to enable wider burst transfers and better bandwidth utilization.
Parallelism Unroll Factors ($TC^{l}_{\text{intra}}$) Unroll factor for each task, determined through tiling, where the intra-tile loop is fully unrolled to expose fine-grained parallelism.
Array Partitioning ($AP_{a,d}$) Number of partitions for array $a$ along dimension $d$ to support parallel accesses.
Compute Padding Padding applied to loop trip counts to increase legal unroll factors.
Code Structure Loop Permutations Legal reordering of loops to optimize data locality and parallelism; synchronized across fused statements.
Tiling Splitting loops into inter- and intra-tile levels for optimization and pipelining.
SLR Assignment ($slr_{t}$) Mapping of each task $t$ to a Super Logic Region (SLR) to enable spatial task distribution.
Design Constraints
Resource Limits Max Array Partitioning Total partitions across all arrays must not exceed architectural limits.
DSP Budget Maximum DSP slices per SLR; used to limit total DSP usage by pipelined/unrolled tasks.
On-Chip Memory Total memory usage by buffered tiles (considering reuse and buffering) must be within BRAM capacity.
Table 2. Design Variables and Architectural Constraints in the Prometheus Optimization Space. This table outlines the key configurable parameters used in Prometheus—including loop tiling, unrolling, array partitioning, and buffer management—along with hardware constraints such as resource limits and memory capacity. The cost model also accounts for concurrent task execution enabled by the dataflow pragma, as well as computation-communication overlap achieved through automatic double or triple buffering. These elements jointly define the valid and efficient design space explored by Prometheus.
Figure 1 illustrates the workflow of Prometheus. The process begins with an input C++ code, from which we extract an intermediate representation containing dependencies, trip counts, schedules, and other relevant information. This extraction is performed using PoCC (Pouchet, 2023), which provides all necessary details for analysis. Next, we construct the task-flow graph based on this extracted information and formulate the corresponding Non-Linear Programming (NLP) problem. Solving this NLP problem determines the theoretical optimal parameters for the design space. Once the NLP solver provides these parameters, we have all the required information to proceed. At this stage, we automatically generate the HLS-C++ code along with all the necessary files to produce the FPGA bitstream, ensuring an efficient and fully automated compilation process.
To support efficient exploration of the design space, Prometheus defines a set of tunable parameters and architectural constraints that capture the key aspects of FPGA acceleration. Table 2 provides a detailed overview of these parameters, including memory-related configurations (e.g., bitwidth, tiling, and reuse levels), parallelism controls (e.g., loop unrolling and array partitioning), and structural transformations (e.g., loop permutations and SLR assignments). In addition, the table outlines hardware constraints such as DSP budgets, on-chip memory limits, and maximum legal array partitioning. These variables and constraints jointly define the feasible design space that the NLP-based optimization engine systematically explores to generate high-performance, hardware-aware FPGA implementations.
The pseudo-code in Listing 4 demonstrates how Prometheus processes the 3mm kernel (Listing 3). Load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFO.
Each taski in the pseudo-code corresponds to the fully unrolled computation of an intra-tile for statement Si in Listing 3. An example of such a task is illustrated in Listing 5.
Arrays $E$, $F$, and $G$ are initialized to zero within their respective tasks (e.g., S0, S2, and S4) and are not preloaded, reducing unnecessary memory overhead.
Prometheus enhances computation efficiency by fusing statements that produce the same outputs (e.g., in Listing 3), and it automatically formulates an NLP problem to determine the theoretically optimal parameters, such as loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding. The framework dynamically adjusts bit widths for arrays like $F$, $D$, and $G$ to achieve a better balance between parallelism and resource usage.
@@ 2.4. Overview of Prometheus
7 task0(E); // S0 + intra-tile loops
@@
10 task1(E, A, B); // S1 + intra-tile loops
@@
18 task2(F); // S2 + intra-tile loops
@@
21 task3(F, C, D); // S3 + intra-tile loops
@@
31 task4(G); // S4 + intra-tile loops
@@
34 task5(G, E, F); // S5 + intra-tile loops
@@
Listing 4: Transformed and Fused 3mm Kernel Pseudocode Generated by Prometheus. This code highlights the result of Prometheus optimizations, including fused task generation, tiled loops, memory buffer management, and pipelined execution with dataflow pragmas.
Efficient computation-communication overlap is achieved through ping-pong buffering, while concurrent task execution and FIFO-triggered dependent tasks maximize overall performance. Fused Tasks 0 and 1 execute concurrently, while Fused Task 2 begins as soon as the data tiles for $F$ and $E$ become available.
To assess the effectiveness of our proposed framework, we evaluate the performance of the 3mm kernel using several state-of-the-art HLS-based FPGA optimization tools. As shown in Table 3, Prometheus significantly outperforms existing frameworks in terms of throughput, achieving 368.36 GF/s. This represents a substantial improvement over Sisyphus (178.97 GF/s) and Stream-HLS (174.00 GF/s), and far exceeds the results of Allo, ScaleHLS, and AutoDSE. The superior performance of Prometheus stems from its holistic design space exploration strategy, which integrates loop transformations, memory tiling, concurrent task execution, and SLR-aware scheduling to deliver optimized and hardware-feasible designs.
Metric Prometheus Sisyphus Stream-HLS Allo ScaleHLS AutoDSE
Throughput (GF/s) 368.36 178.97 174.00 60.40 43.04 1.74
Table 3. Measured Throughput of the 3mm Kernel Using Various FPGA Frameworks. This table presents the runtime throughput (in GF/s) obtained from RTL simulation of the 3mm kernel across multiple frameworks. Prometheus is shown to outperform other methods, including Sisyphus, Stream-HLS, and AutoDSE.
3. Code Transformation
In our dataflow model, we adopt a synchronous dataflow where the sizes of the arrays are known during compile time. This compile-time awareness enables us to construct a precise model that facilitates rigorous optimizations. To leverage FPGA parallelism, we implement an acyclic dataflow graph, ensuring parent nodes do not receive data from their children. While this constraint limits graph configurations and may increase resource usage, it reduces overall latency. To support this structure, we inline (Chen et al., 1993) each function in the input code to generate the required acyclic graph. Our primary objective is to minimize latency within resource constraints by overlapping communication and computation within tasks and executing independent tasks concurrently. To achieve this, we apply various transformations and optimizations, explored in this section, and navigate the design space using an NLP-based approach, detailed in Section 4.
3.1. Dependency Graph Creation
Our process starts with affine C/C++ code as input, which undergoes maximal distribution to ensure each loop body contains only one statement, provided no dependencies exist within the loop. ISCC (Verdoolaege, 2011) verifies the legality of these transformations, ensuring dependencies are preserved. After achieving full distribution, we construct a dependency graph using PoCC (PoCC, [n. d.]). PoCC provides the necessary information about the schedule and dependencies, enabling us to build the graph. In this graph, the nodes represent tasks, while the edges capture data communication arising from inter-task dependencies. Tasks with identical outputs are then merged (when legal), creating fused tasks with output-stationary properties. This ensures that each tile’s output is handled (loaded, computed, and either stored or transmitted) only once. If a dependency prevents distribution, the framework will still function but with a more limited optimization space, making it more effective when distribution is possible.
@@ 3.2. Data-tiling and Padding
In the fused dependency graph, edges represent data communication between tasks as well as between tasks and off-chip memory, involving the transfer of data tiles with specified sizes. Data tiling divides loop iterations into smaller tiles, splitting array accesses into subsets. Each loop $l$ iterating over an array $a$ is divided into an outer loop ($TC^{l}_{inter}$) and an inner loop ($TC^{l}_{intra}$), with feasible permutations applied. For each array and dimension iterated by the loop $l$, the tile factor is a common choice that influences all arrays iterated by this loop.
To optimize memory transfers, padding is applied to arrays in two ways. Simple padding increases the bit width ($BW_{a}$) for efficient transfers while maintaining the original loop trip count. Composite padding adjusts both $BW_{a}$ and the loop trip count ($TC^{l}$) to support unroll factors that do not evenly divide the original trip count, allowing irregular tile sizes and expanding the design space. Tile sizes are consistent within a fused task but vary between tasks. In Listing 4, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
3.3. Fine-grained Parallelism
@@
Data-tile selection involves splitting each loop and permuting them to create two levels of the original loops. Using ISCC (Verdoolaege, 2011), we verify the legality of these permutations, resulting in two loop levels: inter-tile (outer) and intra-tile (inner). If permutation is not feasible, the inter-tile loop retains its legal position. For tasks belonging to the same fused task, we merge their inter-tile loops, which are non-reduction. For instance, in Listing 4, the inter-tile loops on lines 3 and 5 iterate over task0 and task1. The intra-tile loops are fully unrolled, ensuring that all data accesses within the intra-tile remain on-chip. Array partitioning, determined by the unroll factor, ensures data resides in separate BRAM banks for simultaneous access. Reduction loops across inter-tiles are pipelined with an initiation interval ($II=n>1$), where n matches the reduction latency. For instance, in Listing 4, additions take 3 cycles, resulting in $II=3$.
3.4. Loop Order
@@
Due to the full unrolling of the intra-task, there is no need to select the loop order within the intra-tile. We place the inter-tile reduction loops directly above the task and are pipelined. If multiple reduction loops exist, we rank them by the size of their trip counts, placing the loop with the highest trip count innermost to ensure an efficient pipeline. This setup provides the flexibility to choose the order of the inter-tile loops that are not reduction loops. This order will subsequently be determined by the NLP (cf. Section 4). As shown in Listing 4, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 3.
3.5. Automatic Overlapping of Communication and Computation
@@
To overlap communication and computation, we use on-chip buffers. The buffer size and data transfer location are determined by various options explored via NLP (cf. Section 4). Since data must stay on-chip for intra-task computation, transfers occur either below an inter-tile loop or before any loops start. The transfer position determines the data tile size, covering all data accessed below it. To improve data reuse, buffer size can match or exceed the transferred data tile. Similar to transfer location, buffer size is determined by its position relative to inter-tile loops or before any loops. Two boolean variables, $d_{a,l}$ and $t_{a,l}$, define and transfer array $a$ under loop $l$ when set to true. If defined or transferred before loops, $l=0$. Double-buffering is used for read-only or write-only arrays, and triple-buffering for arrays that are both read and written. Following (Cong et al., 2018), we perform an initial load, then overlap loading the next tile with computing the current one.
3.6. Coarse-Grained Parallelism: Automatic Execution of Concurrent Tasks
@@
Due to the use of the dataflow pragma, tasks can begin computation as soon as they have sufficient data, allowing for concurrent execution. This concurrency significantly increases overall performance.
3.7. Memory Transfer
@@
Communication latency is reduced by transferring more data per cycle using increased bit width with padding. Read-only arrays are duplicated in off-chip memory for tasks with multiple reads, eliminating feed-through logic. For non-read-only arrays, data passes between tasks via FIFOs, using the same bit width as for off-chip memory transfer.
4. NLP Formulation
@@
To determine the tile size, loop order, bit width, and memory transfers, we formulate a cost model as a NLP problem aimed at minimizing overall latency. This approach builds upon the methodology proposed in our previous work (Pouget et al., 2025a, 2024, b), which we have further adapted to meet the specific requirements and constraints of our current framework. In our work, we incorporate dataflow considerations along with all the optimizations detailed in Section 3. We employ PoCC (PoCC, [n. d.]) to extract compile-time information such as schedules, loop trip counts, dependencies, and operation counts per statement. ISCC (Verdoolaege, 2011) then generates all legal permutations for each loop body.
Table 4 delineates the sets, variables, and constants utilized in our NLP formulation.
Constants Description
@@
$p_{i}^{l}$ Position of the loop $l$ under the $i$-th permutation
@@
Table 4. Mathematical Notation for Constants, Variables, and Sets in the NLP-Based Optimization Model. This table defines the formal notation used in Prometheus’ Non-Linear Programming (NLP) model for design space exploration, including loop trip counts, bitwidths, SLR mappings, resource limits, and legal transformations.
4.1. Constraints
We now describe the constraints by using the code of Listings 3, 4 and 5.